Commissioning Notes for 20080422 UT

Summary:

LBTtools/IRTC
  • startup and shutdown scripts made.
  • takepic, takeseq, and takeEFpair now properly read IRTC info.
  • Alist and qdisp now work with irtc images (LBTtools tasks), some header caveats.
  • moveOPE now moves in RX and RY, but note below for scale changes.
  • started work on center script to adjust pointing interactively.

irc/irs - several bug fixes and improvements implemented

GCS - unable to move the stages on-axis with presetTelescope command

Some investigations on the origin of the oscillations, inconclusive.

Running Notes/Log (UTtimes):

2:00 Open, waiting for some final SW configuration before starting

3:00 Starting up AGW, waiting for UMAC temp to reach >10C
  • We should make a startup script and make agwuser a sudoer so that the script can spawn the first xterm and execute the check4nport command. Once the OK is reached, a second script can spawn the second xterm and execute the oacserver command.

03:33 pointto WT10_256, star appeared at bottom of detector! Offset 3.0 -12.0 to center

03:40 persistent irs socket problem on GetImage, rebooting IRTC control computer...

03:50 One image taken, subsequent calls still have socket errors. Rebooting again.
  • Still persistent socket problems.

Note: movie mode shows a systematic, cyclic oscillation at 45 degrees to NE (is this in AZ or EL?). See more discussion below.

06:34 After some debugging we are going on with GCS checkout. AGW started and homed.

WT10_294:

06:45 First "GUIDE" mode failed as expected because GCS was not running

06:50 Next GUIDE failed, problem on TCS side...working on it.

WT10_298: at PA=180 to stay off rotator limits

09:30 moveOPE's RX and RY parameters now work, but there are scaling issues:
  • For M1 - an RX or RY request shifts the image on the IRTC by double that amount.
  • For M2 - an RX or RY request shifts the image on the IRTC by 2.0/9.6 that amount.
  • Factor of 9.6 is approximate.

09:48 Testing GCS again. Failed for some communication between GCS and PCS.

'IssueTrak' 1292 - AGW's UMAC went up to a max of ~29C and has since come back down to a stable ~28.2C.

'IssueTrak' 1293 - agw-control computer's time set to CEST, probably has not affected operations.

Oscillations with a 2s period seen in movie mode, where the PSF is stationary for about 1s and then oscillates for ~1s. Michele confirmed that the SkyPatchUpdater runs at this speed, so we investigated changing the SkyPatchUpdater to 4s and 1s periods. No change was seen on the movie. We also checked varying the exposure time on the IRTC (with the frame rate set to 100,000), at 1ms, 10ms 15ms no significant changes were seen in the oscillatory behavior. It does not look like it is the SkyPatchUpdater, but this bears more investigation. What could it be?

JMH: It looks like this could be the same oscillation that Doug reported on Saturday (Issue #1284) - except that Doug was looking at it (via data cubes) at higher speed. It looks to my eye that there is a beat frequency of a 3-4 Hz oscillation so that the amplitude is varying from small to ~1 arcsec on a timescale of ~3 sec. Chris and I turned the M1 Cell outer loop servos on and off several times. It seems that the outer loop servos are not the source of the oscillation.

Note from Michele: Please note that I modified the PCSInstrument.conf file located in /home/telescope/TCS/Configuration/PCS

to have a LEFTZEROPOINT for IRTCBGF (sorry software and engineering have slightly different ways of referencing the focal stations) of 119.1 degrees. This is 135.0 + 64.1. The "pre-modified by Michele file" is 21AprPCSInstrument.conf.

12:00 Closed, twilight.

-- DavidThompson - 22 Apr 2008
Topic revision: r7 - 22 Apr 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback